OpenStudy (anonymous):

Hey Smarties, would you guys be kind and help me find the inverse for the function: f(x)=(x^2)/3

OpenStudy (jdoe0001):

to get the inverse of a function just switch the letters about y=(x^2)/3 inverse will be x=(y^2)/3

OpenStudy (marissalovescats):

For an aninverse function, use y= and switch the place of x=y.

OpenStudy (jdoe0001):

then you can just solve by "y"

OpenStudy (marissalovescats):

So what @jdoe0001 Just showed :)

OpenStudy (jdoe0001):

meow!

OpenStudy (anonymous):

jdoe is wrong its 5